# Nan-Yow Chen4, Samuel Yen-Chi Chen81, En-Jui Kuo92, Hsi-Sheng Goan67103

### v1

@@ I Introduction
    We achieve a substantial 99.5% reduction in trainable parameters compared to classical LSTMs while maintaining or improving predictive performance across multiple datasets.
@@ III-A2 Integration of Quantum-inspired Kolmogorov–Arnold Networks into LSTM
where $M$ is a fixed Hermitian observable (e.g., $\sigma_{z}$). The corresponding data re-uploading circuit $U(u;\boldsymbol{\theta})$ consists of a sequence of $L$ parameterized quantum blocks:
$U(u;\boldsymbol{\theta})=W^{(\ell+1)}\prod_{\ell=1}^{L}\Big[W^{(\ell)}\;\exp\!\big(-i\,\tfrac{a^{(\ell)}u+b^{(\ell)}}{2}\,H\big)\ \Big].$ (4)
Here $H$ denotes a fixed Hermitian generator, $a^{(\ell)}$ and $b^{(\ell)}$ are scalar encoding parameters, and $W^{(\ell)}(\boldsymbol{\theta})$ represents a trainable single-qubit variational block. Stacking $L$ such re-uploading layers with the help of $a^{(\ell)}$ and $b^{(\ell)}$ endows the activation $\phi_{g,p}$ with an exponentially enriched Fourier spectrum, enabling compact yet expressive nonlinear representations.
@@ IV-C Result Analysis
V Conclusion
This study has demonstrated that incorporating the QKAN framework into LSTM architectures substantially enhances temporal sequence modeling capabilities across both synthetic and real-world datasets. The proposed QKAN-LSTM and HQKAN-LSTM models utilize the expressive power of quantum mappings to enrich classical recurrent representations, creating a compact yet powerful feature encoding for temporal dependencies. This is achieved by enhancing frequency adaptability and enabling an exponentially enriched spectral representation without requiring complex multi-qubit entanglement. This integration allows the models to deliver higher predictive accuracy and faster convergence while requiring far fewer trainable parameters and quantum resources, fundamentally boosting computational efficiency and scalability.
@@ Acknowledgment
Y.C. Hsu, J.-C. Jiang, C.-H. Lin and K.C. Peng thank the National Center for High-Performance Computing (NCHC), National Institutes of Applied Research (NIAR), Taiwan, for providing computational and storage resources supported by National Science and Technology Council (NSTC), Taiwan, under Grants No. NSTC 114-2119-M-007-013. H.-S. Goan acknowledges support from the NSTC, Taiwan, under Grants No. NSTC 113-2112-M-002-022-MY3, No. NSTC 113-2119-M-002-021, No. 114-2119-M-002-018, No. NSTC 114-2119-M-002-017-MY3, from the US Air Force Office of Scientific Research under Award Number FA2386-23-1-4052 and from the National Taiwan University under Grants No. NTU-CC-114L8950, No. NTU-CC114L895004 and No. NTU-CC-114L8517. H.-S. Goan is also grateful for the support of the “Center for Advanced Computing and Imaging in Biomedicine (NTU-114L900702)” through the Featured Areas Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education (MOE), Taiwan, the support of Taiwan Semiconductor Research Institute (TSRI) through the Joint Developed Project (JDP) and the support from the Physics Division, National Center for Theoretical Sciences, Taiwan.

### v2

# arXiv:2512.05049v2[quant-ph] 27 Feb 2026

@@ I Introduction
    We achieve a substantial 79% reduction in trainable parameters compared to classical LSTMs while maintaining or improving predictive performance across multiple datasets.
@@ III-A2 Integration of Quantum-inspired Kolmogorov–Arnold Networks into LSTM
where $M$ is a fixed Hermitian observable (e.g., $\sigma_{z}$). The corresponding data re-uploading circuit $U(u;\boldsymbol{\theta})$ consists of a sequence of $r$ parameterized quantum blocks:
$U(u;\boldsymbol{\theta})=W^{(r+1)}\prod_{\ell=r}^{1}\Big[\exp\!\big(-i\,\tfrac{a^{(\ell)}u+b^{(\ell)}}{2}\,G\big)W^{(\ell)}\ \Big].$ (4)
Here $G$ denotes a fixed Hermitian generator, $a^{(\ell)}$ and $b^{(\ell)}$ are scalar encoding parameters, and $W^{(\ell)}(\boldsymbol{\theta})$ represents a trainable single-qubit variational block. Stacking $r$ such re-uploading layers with the help of $a^{(\ell)}$ and $b^{(\ell)}$ endows the activation $\phi_{g,p}$ with an exponentially enriched Fourier spectrum, enabling compact yet expressive nonlinear representations.
@@ IV-C Result Analysis
V Discussion
In this work, the performance gains over the classical baseline are consistently associated with replacing the gate-wise affine maps in LSTM by DARUAN-based QVAFs. DARUAN implements a single-qubit data re-uploading circuit with trainable pre-processing weights, yielding a compact and highly expressive univariate nonlinearity whose controllable Fourier spectrum can expand exponentially with the number of re-uploading repetitions [29]. When deployed inside recurrent gates, this spectral adaptability is particularly beneficial for time series exhibiting mixed periodicity, burstiness, and multi-scale temporal patterns, as the gate transformations can represent richer nonlinear filtering behavior without increasing hidden width.
Beyond expressivity, integrating QKAN into LSTM provides a structural advantage inherited from KAN-style networks: each gate transformation is realized as an additive composition of learnable univariate functions. This decomposition offers a more transparent mechanism for analyzing how individual input channels and hidden-state components influence the forget/input/output decisions, enabling gate-level interpretability and more targeted diagnostics of the learned temporal control policy. Compared to monolithic dense transformations, such edge-wise QVAF parameterization also encourages parameter reuse across time steps and alleviates redundancy in gate modeling.
From an optimization perspective, QKAN activations are computed as bounded expectation values of single-qubit observables, which naturally constrains gate pre-activations and acts as an implicit regularizer. In recurrent learning, where backpropagation through time is sensitive to unstable gate dynamics, these bounded nonlinearities complement the outer sigmoid/tanh nonlinearities to improve training stability. Moreover, because QKAN avoids multi-qubit entanglement by construction, it reduces dependence on deep entangling circuit structures and the associated hardware and optimization burdens common in VQC-based recurrent models [13]. In particular, scaling multi-qubit VQCs often amplifies barren plateau phenomena [64] and exacerbates the impact of imperfect two-qubit gates [65, 66] and shot-noise-dominated measurements [67, 68, 69, 70, 71]. By contrast, the single-qubit and additive structure of QKAN mitigates these issues and supports progressive expressivity scaling via the layer extension strategy (increasing the re-uploading depth to enrich frequency content while preserving previously learned features) [29], which is well aligned with stable optimization in recurrent settings.
Finally, QKAN-based gates can improve practical efficiency. Since each DARUAN operates on a two-level quantum state, the resulting computations are lightweight and can be implemented with GPU-friendly tensor operations, enabling efficient classical simulation and deployment while maintaining quantum-inspired expressivity [29]. In our experiments, this translates into a favorable accuracy–efficiency trade-off: QKAN-LSTM and its autoencoder-like HQKAN-LSTM variant achieve competitive or superior forecasting performance while substantially reducing the number of trainable parameters relative to standard LSTM and requiring fewer quantum resources than multi-qubit QLSTM baselines. Furthermore, ref. [29] shows that HQKAN can serve as a drop-in replacement for MLP layers with a substantially reduced parameter count and lower training memory, while maintaining comparable training throughput. In our setting, the HQKAN-LSTM inherits this efficiency advantage by compressing the latent dimensionality of the gate transformations, thereby reducing the overall parameterization and the corresponding memory footprint relative to a standard LSTM.
VI Conclusion
This study has demonstrated that incorporating the QKAN framework into LSTM architectures substantially enhances temporal sequence modeling capabilities across both synthetic and real-world datasets. The proposed QKAN-LSTM and HQKAN-LSTM models utilize the expressive power of quantum mappings to enrich classical recurrent representations, creating a compact yet powerful feature encoding for temporal dependencies. This is achieved by enhancing frequency adaptability and enabling an exponentially enriched spectral representation without requiring complex multi-qubit entanglement. This integration allows the models to deliver higher predictive accuracy and faster convergence while requiring significantly fewer trainable parameters and quantum resources, fundamentally boosting computational efficiency and scalability.
@@ Acknowledgment
Y.C. Hsu, J.-C. Jiang, C.-H. Lin and K.C. Peng thank the National Center for High-Performance Computing (NCHC), National Institutes of Applied Research (NIAR), Taiwan, for providing computational and storage resources supported by National Science and Technology Council (NSTC), Taiwan, under Grants No. NSTC 114-2119-M-007-013. H.-S. Goan acknowledges support from the NSTC, Taiwan, under Grants No. NSTC 113-2112-M-002-022-MY3, No. NSTC 113-2119-M-002-021, No. NSTC 114-2119-M-002-018, No. NSTC 114-2119-M-002-017-MY3, and from the National Taiwan University under Grants No. NTU-CC-115L8937, No. NTU-CC115L893704 and No. NTU-CC-115L8512. H.-S. Goan is also grateful for the support of the “Center for Advanced Computing and Imaging in Biomedicine (NTU-115L900702)” through the Featured Areas Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education (MOE), Taiwan, the support of Taiwan Semiconductor Research Institute (TSRI) through the Joint Developed Project (JDP) and the support from the Physics Division, National Center for Theoretical Sciences, Taiwan.